OpenStudy (anonymous):

Find the LCM of 32 and 27. A. 288 B. 864 C. 432 D. 216 @some.random.cool.kid@geerky42@Bee_see

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

@Bee_see

OpenStudy (anonymous):

@some.random.cool.kid

geerky42 (geerky42):

\(32 = 2^5\) and \(27 = 3^3\) So \(\text{GCF}(32,27) = 1\) So that mean to find \(\text{LCM}\), you just multiply \(32\) and \(27\).

OpenStudy (anonymous):

What is the GCF of 36 and 84? A. 4 B. 6 C. 12 D. 18

geerky42 (geerky42):

\[36 = 2\times2\times3\times3\\ 84 = 2\times2\times3\times7\]

geerky42 (geerky42):

\[36 = 12\times3\\ 84 = 12\times7\]

geerky42 (geerky42):

So \(\text{GCF}(36,84)=\boxed{12}\)